# TagAlign: Improving Vision-Language Alignment with Multi-Tag Classification

### v1

@@ Abstract
The crux of learning vision-language models is to extract semantically aligned information from visual and linguistic data. Existing attempts usually face the problem of coarse alignment, e.g., the vision encoder struggles in localizing an attribute-specified object. In this work, we propose an embarrassingly simple approach to better align image and text features with no need of additional data formats other than image-text pairs. Concretely, given an image and its paired text, we manage to parse objects (e.g., cat) and attributes (e.g., black) from the description, which are highly likely to exist in the image. It is noteworthy that the parsing pipeline is fully automatic and thus enjoys good scalability. With these parsed semantics as supervision signals, we can complement the commonly used image-text contrastive loss with the multi-tag classification loss. Extensive experimental results on a broad suite of semantic segmentation datasets substantiate the average 3.65% improvement of our framework over existing alternatives. Furthermore, the visualization results indicate that attribute supervision makes vision-language models accurately localize attribute-specified objects. Project page can be found at https://qinying-liu.github.io/Tag-Align/
$*{}^{*}$ Intern at Ant Group.      ${}^{\dagger}$ Corresponding authors.
@@ 1 Introduction
Figure 1: Illustration on various tag supervisions (e.g., object and attribute) for open-vocabulary semantic segmentation. The vanilla CLIP struggles in localizing an attribute-specified object. When introducing object supervision (as depicted in the third column), CLIP can focus on the more accurate region of text-specified object (e.g., cat). In addition, adding attribute supervision brings CLIP a stronger understanding of visual attribute-related concepts (as depicted in the fourth column).
Key aspect in the process of learning vision-language (V&L) encoders, e.g., CLIP [27], is to effectively extract semantically aligned information from both visual and linguistic data. A well-aligned vision-language encoder stands out for its remarkable versatility, which includes unprecedented quality in text-to-image generation, as well as the improvement in multi-modality understanding tasks. However, existing attempts usually face the problem of coarse alignment, e.g., the vision encoder struggles in localizing an attribute-specified object (as depicted in Fig. 1a), making it hard to achieve precise localization of arbitrary concepts.
@@
Thereupon, recent researchers have turned their attention to how to automatically parse the data of image-text pairs, especially text. Concretely, automatic text parsing tool (i.e., NLTK) has become a de-facto approach in V&L field, which can generate some tags from caption of image. While these tags can bring performance improvements in V&L models, these endeavors, incorporating much noise from NLTK, may restrict the models’ ability to reach the upper limit of accurately perceiving semantic concepts. For collecting higher-quality data, Tag2Text [12] designs a universal and unified label system with an efficient data annotation engine, which generate about $4,000$ tags from some image-text datasets (e.g., CC12M). Based on these tags, the series works of RAM [12, 45] explore an image-tag recognition system with decoder module, which can classify commonly human-used categories. Additionally, DiHT [26] distills the knowledge from an extra multi-tag classification model, trained on collected private tags, only into the visual encoder of CLIP. These methods focus on improving the classification performance of visual encoders, without explicitly exploring the alignment between vision and language. A core viewpoint still worth exploring: how to utilize various multi-tags to help alignment between the image and text embeddings from V&L encoders? I argue that there are two important questions in the training design of multi-modal encoders based on multi-tags. The first one is the impact about the number of tag. Even using high-end tag signals in prior methods, a limited number of tags makes them fail to fully unleash the potential of multiple tags. Furthermore, the second one is that there is a lack of comprehensive exploration on how to effectively utilize various tags (e.g., attribute and objects).
@@
In this paper, we propose an embarrassingly simple approach to improve the alignment between image and text features without the need for additional data formats beyond image-text pairs. Specifically, given an image and its paired text, we automatically parse objects (e.g., cat) and attributes (e.g., brown) from the description, which are highly likely to exist in the image. This parsing pipeline is implemented by a large language model, which is high-end and fully automatic, ensuring quality and scalability. By using these parsed semantics as supervision signals, we can complement the commonly used image-text contrastive loss with a multi-tag classification loss, enabling the model to better capture the visual semantic concepts referred to in the texts. Through that, our method becomes capable of accurately localizing text-specified objects (e.g., attribute-specified object), fulfilling more precise alignment between image and text. Extensive experimental results as shown in Fig. 2 on a broad suite of semantic segmentation datasets (10 datasets) substantiate the average 3.65% improvement of our framework over existing alternatives. Moreover, it is mentioned that we only use the CC12M dataset with the parsed text as the training set, indicating that our method is both efficient and effective.
@@
Besides remarkable performance on open-vocabulary semantic segmentation tasks, TaskAlign also enjoys the high flexibility to incorporate attribute information. Concretely, the visualization results, as shown in Fig. 1, show that the inclusion of attribute supervision enables CLIP to accurately localize the attribute-specified object (‘brown cat’). This finding indicates that different types of tag supervision have the potential to enhance the ability of cross-modal alignment in different aspects of CLIP.
Figure 2: Comparison on various semantic segmentation benchmarks. TagAlign surpasses existing methods with a large margin on most benchmarks.
@@ 2 Related Work
Figure 3: The framework of TagAlign. At the core of TagAlign are two key components: a) LLM-aided tag parsing that parses the image captions into diverse tags (i.e., objects and attributes); b) Multi-tag classification that utilizes the parsed tags to supervise the model training.
@@ 3 Method
Our method comprises two key ingredients: 1) LLM-aided tag parsing, which aims to extract diverse tags (e.g., objects and attributes) from the captions of the images; 2) multi-tag classification, which targets to guide the model training via a multi-tag classification objective. To be specific, given an image-caption pair, we prompt LLM to analyze the caption and extract tags from it. As a result, a large tag list can be established from the whole image-caption dataset. By indexing on this tag list, we transform the tags of each image to a one-hot label, which enables us to use the label to supervise the model training with a simple multi-tag classification loss.
@@ 3.1 LLM-Aided Tag Parsing
In this section, we provide a detailed explanation of our proposed method for extracting tags from captions. Traditional methods [40, 41] commonly employ the Natural Language ToolKit (NLTK) [21] for text parsing. However, as illustrated in Fig. 6, NLTK analyzes the texts solely based on Part-of-Speech (POS), leading to challenges in understanding the scene described in the caption. On the other hand, large language models (LLMs) have demonstrated exceptional instruction-following capabilities, rendering them more controllable and flexible in parsing the texts. In light of this, we propose to prompt existing LLM to analyze and extract tags (i.e., objects and attributes) from given captions. Without loss of generality, we focus on objects for illustration purposes in the following.
Prompt Engineering. The prompts we employ consist of specific rules and examples. Firstly, a set of rules is crafted to constrain the LLM’s responses, ensuring a focus on tangible and visible objects likely to exist in the image. Secondly, examples are provided to guide the LLM in generating the desired response format. These examples leverage the LLM’s in-context learning ability to instruct it on how to structure and format the output accurately. We refer to Supplementary Material for more details about the prompts that we have used.
With the prompts in place, we query the LLM to extract tags from the captions. In Fig. 6, we offer examples to illustrate the comparison between tags extracted by the LLM and NLTK. As demonstrated, the LLM offers two significant advantages over NLTK. Firstly, it demonstrates a more accurate understanding of the scene described in the caption, resulting in a more precise listing of tangible and visible objects. Secondly, the LLM performs word splitting in a more logical manner, mitigating the issue of segmenting entire objects into multiple components.
Tag List Construction. Given an image-caption dataset consisting of image-caption pairs, each caption is sequentially processed by the LLM to extract tags. Subsequently, a comprehensive tag list is generated by aggregating all extracted tags and eliminating duplicates. To ease the pre-training process, we selectively retain only the top $K$ tags that appear most frequently. The most frequent object and attribute tags extracted by LLM are illustrated in Fig. 4.
@@
Figure 4: Illustration of the most frequent tags in the object tag list (a) and attribute tag list (b). The font size corresponds to the tag’s frequency of occurrence in the tag list: the more frequent tag uses a larger font size.
@@ 3.2 Multi-Tag Classification
As shown in Fig. 3, for an image ${\bm{x}}$, we first pass it through the pre-trained image encoder of CLIP to obtain the patch embeddings (excluding the [CLS] embedding). These patch embeddings are then processed by a lightweight projector, resulting in the final patch feature map denoted as ${\mathbf{Z}}\in\mathbb{R}^{L\times D}$, where $L$ is the number of patches. With these patch-level embeddings, our objective is to achieve fine-grained image-text alignment, despite the availability of only image-level labels for training.
To this end, we formulate the problem as a special case of Multiple Instance Learning (MIL), in which a “bag” (image) contains multiple “instances” (patches), and the labels of these instances collectively determine the label of the bag. Following the standard practice in MIL [37], it is necessary to train a tag classifier to predict the bag-level label. Once trained, the tag classifier can be applied to the instances (patches), where the instances triggering the bag label should exhibit high activations.
To implement this approach, we first apply average pooling over the patch feature map to obtain the bag(image)-level representation ${\bm{z}}\in\mathbb{R}^{D}$, calculated as ${\bm{z}}=\frac{1}{L}\sum_{l=1}^{L}{\mathbf{Z}}_{l}$. Subsequently, ${\bm{z}}$ is fed into a tag classifier to predict tag classification scores. To leverage CLIP’s zero-shot transfer capability, instead of using randomly initialized weights, we leverage the pre-trained text encoder from CLIP to generate fixed weights for the $K$ tags. More specifically, the tags are firstly prompted with the template “A photo of a {tag}”, and then the prompted texts are processed by the text encoder, resulting in tag embeddings, denoted by $\{{\bm{c}}_{1},..,{\bm{c}}_{K}\}$. These tag embeddings serve as the weights for the tag classifier. As a result, the tag classification scores can be computed as the relative similarity between the bag-level representation and the tag embeddings. Formally, the $k$-th tag classification score ${\bm{p}}_{k}$ is expressed as:
@@
where $\rho$ is a learnable temperature factor. Finally, we optimize the model through a multi-label cross-entropy loss between the tag classification scores ${\bm{p}}$ and the binarized tag labels ${\bm{y}}$:
@@
Figure 5: Statistic of object & attribute tags from LLM parser.
@@
This equation is similar to the noise-contrastive estimation (NCE) [9, 10] process. Minimizing Eq. 2 is thereby tantamount to maximizing the lower bound of mutual information (MI) between the textual tag embeddings ${\bm{c}}$ and the visual image embedding ${\bm{z}}$. More specifically, through the optimization of Eq. 2, the model is trained to align visual and text representations. This alignment is achieved by encouraging positive textual tag embeddings to be close to the visual image embedding and simultaneously ensuring that negative textual tag embeddings are farther away from the visual image embedding. As the image embedding represents a “bag” of snippet embeddings in MIL, this alignment facilitates capturing the fine-grained semantic relationship between patches and tags.
@@
Handing Long-tailed Distribution. As depicted in Fig. 5, the distribution of the tags is highly imbalanced, even after filtering certain tags in the tail. To mitigate this issue, we propose to utilize the balanced softmax strategy [29]. Specifically, we adjust prediction logits by multiplying by the tag frequencies. That is, we transform Eq. 1 into
@@ 3.3 Training and Inference
Here, with a slight abuse of notation, $N$ represents the number of image-text pairs in a batch, and ${{\bm{z}}}_{n}$ and ${{\bm{t}}}_{n}$ denote the visual embedding of the $n$-th image and $n$-th caption within the batch.
@@
Inference. To generate the final segmentation map, we first reshape the patch feature map ${\mathbf{Z}}$ to spatial shape: ${\mathbf{Z}}\in\mathbb{R}^{L\times D}\rightarrow{\mathbf{Z}}\in\mathbb{R}^{H\times W\times D}$, where $L=HW$. Next, we compute the similarity map between the patch features and the class embeddings. To generate class embeddings, we use the prompt template “a photo of a {label}”. This similarity map captures the degree of similarity between each patch and the different classes. Afterward, we upsample the similarity map to match the original image size through linear interpolation. Finally, we apply the $\operatorname{Argmax}$ function along the class dimension of the similarity map, thereby assigning each pixel to a particular class.

### v3

@@ Abstract
The crux of learning vision-language models is to extract semantically aligned information from visual and linguistic data. Existing attempts usually face the problem of coarse alignment, e.g., the vision encoder struggles in localizing an attribute-specified object. In this work, we propose an embarrassingly simple approach to better align image and text features with no need of additional data formats other than image-text pairs. Concretely, given an image and its paired text, we manage to parse objects (e.g., cat) and attributes (e.g., black) from the description, which are highly likely to exist in the image. It is noteworthy that the parsing pipeline is fully automatic and thus enjoys good scalability. With these parsed semantics as supervision signals, we can complement the commonly used image-text contrastive loss with the multi-tag classification loss. Extensive experimental results on a broad suite of semantic segmentation datasets substantiate the average 5.2% improvement of our framework over existing alternatives. Furthermore, the visualization results indicate that attribute supervision makes vision-language models accurately localize attribute-specified objects. Project page can be found at https://qinying-liu.github.io/Tag-Align.
Equal contribution   🖂Corresponding authors   $\dagger$ Interns at Ant Group
@@ 1 Introduction
Figure 1: (a)Illustration of the effect of various tag supervisions (e.g., object and attribute) for open-vocabulary semantic segmentation. The vanilla CLIP struggles in localizing an attribute-specified object. When introducing object supervision (as depicted in the third column), CLIP can focus on the more accurate region of text-specified object (e.g., cat). In addition, adding attribute supervision brings CLIP a stronger understanding of visual attribute-related concepts (as depicted in the fourth column). (b)Performance Comparison on various semantic segmentation benchmarks. TagAlign outperforms existing methods by a large margin on all these benchmarks. Best viewed in color.
Key aspect in the process of learning vision-language (V&L) encoders, e.g., CLIP [28], is to effectively extract semantically aligned information from both visual and linguistic data. A well-aligned vision-language encoder stands out for its remarkable versatility, which includes unprecedented quality in text-to-image generation, as well as the improvement in multi-modality understanding tasks. However, existing attempts usually face the problem of coarse alignment, e.g., the vision encoder struggles in localizing an attribute-specified object (as depicted in Fig. 1(a), making it hard to achieve precise localization of arbitrary concepts.
@@
Thereupon, recent researchers have turned their attention to how to automatically parse the data of image-text pairs, especially text. Concretely, automatic text parsing tool (i.e., NLTK) has become a de-facto approach in V&L field, which can generate some tags from caption of image. While these tags can bring performance improvements in V&L models, these endeavors, incorporating much noise from NLTK, may restrict the models’ ability to reach the upper limit of accurately perceiving semantic concepts. For collecting higher-quality data, Tag2Text [13] designs a universal and unified label system with an efficient data annotation engine, which generate about $4,000$ tags from some image-text datasets (e.g., CC12M). Based on these tags, the series works of RAM [13, 46] explore an image-tag recognition system with decoder module, which can classify commonly human-used categories. Additionally, DiHT [27] distills the knowledge from an extra multi-tag classification model, trained on collected private tags, only into the visual encoder of CLIP. These methods focus on improving the classification performance of visual encoders, without explicitly exploring the alignment between vision and language. A core viewpoint still worth exploring: how to utilize various multi-tags to help alignment between the image and text embeddings from V&\&&L encoders? I argue that there are two important questions in the training design of multi-modal encoders based on multi-tags. The first one is the impact about the number of tag. Even using high-end tag signals in prior methods, a limited number of tags makes them fail to fully unleash the potential of multiple tags. Furthermore, the second one is that there is a lack of comprehensive exploration on how to effectively utilize various tags (e.g., attribute and objects).
@@
In this paper, we propose an embarrassingly simple approach to improve the alignment between image and text features without the need for additional data formats beyond image-text pairs. Specifically, given an image and its paired text, we automatically parse objects (e.g., cat) and attributes (e.g., brown) from the description, which are highly likely to exist in the image. This parsing pipeline is implemented by a large language model, which is high-end and fully automatic, ensuring quality and scalability. By using these parsed semantics as supervision signals, we can complement the commonly used image-text contrastive loss with a multi-tag classification loss, enabling the model to better capture the visual semantic concepts referred to in the texts. Through that, our method becomes capable of accurately localizing text-specified objects (e.g., attribute-specified object), fulfilling more precise alignment between image and text. Extensive experimental results as shown in Fig. 1(b) on a broad suite of semantic segmentation datasets (10 datasets) substantiate the average 5.2% improvement of our framework over existing alternatives. Moreover, it is mentioned that we only use the CC12M dataset with the parsed text as the training set, indicating that our method is both efficient and effective.
@@
Besides remarkable performance on open-vocabulary semantic segmentation tasks, TaskAlign also enjoys the high flexibility to incorporate attribute information. Concretely, the visualization results, as shown in Fig. 1(a), show that the inclusion of attribute supervision enables CLIP to accurately localize the attribute-specified object (‘brown cat’). This finding indicates that different types of tag supervision have the potential to enhance the ability of cross-modal alignment in different aspects of CLIP.
@@ 2 Related Work
Figure 2: The framework of TagAlign. At the core of TagAlign are two key components: a) LLM-aided tag parsing that parses the image captions into diverse tags (i.e., objects and attributes); b) Multi-tag classification that utilizes the parsed tags to supervise the model training. Best viewed in color.
@@ 3 Method
Our method comprises two key ingredients: 1) LLM-aided tag parsing, which aims to extract diverse tags (e.g., objects and attributes) from the captions of the images; 2) multi-tag classification, which targets to guide the model training via a multi-tag classification objective. To be specific, given an image-caption pair, we prompt LLM to analyze the caption and extract tags from it. As a result, a large tag list can be established from the whole image-caption dataset. By indexing on this tag list, we transform the tags of each image to a binarized label, which enables us to use the label to supervise the model training with a simple multi-tag classification loss.
@@ 3.1 LLM-Aided Tag Parsing
In this section, we provide a detailed explanation of our proposed method for extracting tags from captions. Traditional methods [41, 42] commonly employ the Natural Language ToolKit (NLTK) [22] for text parsing. However, as illustrated in Fig. 4, NLTK analyzes the texts solely based on Part-of-Speech (POS), leading to challenges in understanding the scene described in the caption. On the other hand, large language models (LLMs) have demonstrated exceptional instruction-following capabilities, rendering them more controllable and flexible when parsing the texts. In light of this, we propose to prompt existing LLM to analyze and extract tags (i.e., objects and attributes) from given captions. Without loss of generality, we focus on objects for illustration purposes in the following.
Prompt Engineering. The prompts we employ consist of specific rules and examples. Firstly, a set of rules is crafted to constrain the LLM’s responses, ensuring a focus on tangible and visible objects likely to exist in the image. Secondly, examples are provided to guide the LLM in generating desired response format. These examples leverage the in-context learning ability of LLM to instruct it on how to structure and format the output accurately. The Vicuna-33b [5] is used as the LLM to extract tags in our implementation. We refer to Appendix for more details about the used prompts.
With the prompts in place, we query LLM to extract tags from the captions. In Fig. 4, we offer examples to illustrate the comparison between tags extracted by LLM and NLTK. As demonstrated, LLM offers two significant advantages over NLTK. Firstly, it demonstrates a more accurate understanding of the scene described in the caption, resulting in a more precise listing of tangible and visible objects. Secondly, LLM performs word splitting in a more logical manner, mitigating the issue of segmenting entire objects into multiple components.
Tag List Construction. Given an image-caption dataset consisting of image-caption pairs, each caption is sequentially processed by the LLM to extract tags. Subsequently, a comprehensive tag list is generated by aggregating all extracted tags and eliminating duplicates. To ease the pre-training process, we selectively retain only the top $K$ tags that appear most frequently. The most frequent object and attribute tags extracted by LLM are illustrated in Fig. 3(a)(b).
@@
Figure 3: Illustration of the most frequent tags in the (a) object tag list and (b) attribute tag list. The font size corresponds to the tag’s frequency of occurrence in the tag list: the more frequent tag uses a larger font size. (c) Statistic of object & attribute tags from LLM parser. The distribution of the tags exhibits a long-tailed pattern. Best viewed in color.
@@ 3.2 Multi-Tag Classification
As shown in Fig. 2, for an image ${\bm{x}}$, we first pass it through the pre-trained image encoder of CLIP to obtain the patch embeddings (excluding the [CLS] embedding). These patch embeddings are then processed by a lightweight projector, resulting in the final patch feature map denoted as ${\mathbf{Z}}\in\mathbb{R}^{L\times D}$, where $L$ is the number of patches. With these patch-level embeddings, our objective is to achieve fine-grained image-text alignment, despite the availability of only image-level labels for training. This is a special case of Multiple Instance Learning (MIL), in which a “bag” (image) contains multiple “instances” (patches), and the labels of these instances collectively determine the label of the bag. Following the standard practice in MIL [38], it is necessary to train a classifier for predicting the bag-level labels (i.e., image-level tags). Once trained, the classifier can be applied to the instances (patches), where the instances triggering the bag label should exhibit high activations.
To implement this approach, we first apply average pooling over the patch feature map to obtain the image-level representation ${\bm{z}}\in\mathbb{R}^{D}$, calculated as ${\bm{z}}=\frac{1}{L}\sum_{l=1}^{L}{\mathbf{Z}}_{l}$. Subsequently, ${\bm{z}}$ is fed into a tag classifier to predict tag classification scores. To leverage the zero-shot transfer capability of CLIP, instead of using randomly initialized weights, we leverage the pre-trained text encoder from CLIP to generate fixed weights for the $K$ tags. More specifically, the tags are firstly prompted with the template “A photo of a {tag}”, and then the prompted texts are processed by the text encoder, resulting in tag embeddings, denoted by $\{{\bm{c}}_{1},..,{\bm{c}}_{K}\}$. These tag embeddings serve as the weights for the tag classifier. As a result, the tag classification scores can be computed as the relative similarity between the image-level representation and the tag embeddings. Formally, the $k$-th tag classification score ${\bm{p}}_{k}$ is expressed as:
@@
where $\rho$ is a learnable temperature scalar. Finally, we optimize the model through a multi-label cross-entropy loss between the tag classification scores ${\bm{p}}$ and the binarized tag labels ${\bm{y}}$:
@@
This equation is similar to the noise-contrastive estimation (NCE) [10, 11] process. Minimizing Eq. 2 is thereby tantamount to maximizing the lower bound of mutual information (MI) between the textual tag embeddings ${\bm{c}}$ and the visual image embedding ${\bm{z}}$. More specifically, through the optimization of Eq. 2, the model is trained to align visual and text representations. This alignment is achieved by encouraging positive textual tag embeddings to be close to the visual image embedding and simultaneously ensuring that negative textual tag embeddings are farther away from the visual image embedding. As the image embedding represents a “bag” of patch embeddings in MIL, this alignment facilitates capturing the fine-grained semantic relationship between patches and tags.
@@
Handing Long-tailed Distribution. As depicted in Fig. 3(c), the distribution of the tags is highly imbalanced, even after filtering certain tags in the tail. To mitigate this issue, we propose to utilize the balanced softmax strategy [30]. Specifically, we adjust prediction logits by multiplying by the tag frequencies. That is, we transform Eq. 1 into
@@ 3.3 Training and Inference
Here, with a slight abuse of notation, $N$ represents the number of image-text pairs within a batch, and ${{\bm{z}}}_{n}$ and ${{\bm{t}}}_{n}$ denote the feature of the $n$-th image and the feature of the $n$-th caption, respectively.
@@
Inference. We quantitatively evaluate our model over segmentation datasets to showcase its ability in achieving fine-grained image-text alignment. For generating the segmentation results, we begin by reshaping the patch feature map ${\mathbf{Z}}\in\mathbb{R}^{L\times D}$ to a spatially structured format ${\mathbf{Z}}\in\mathbb{R}^{H\times W\times D}$, where $L=HW$. Next, we compute the similarity map between the patch features and the class embeddings. The class embeddings are generated by applying the prompt template “a photo of a {class name}” to the class names. Afterward, we upsample the similarity map to match the original image size through linear interpolation. Finally, the $\operatorname{Argmax}$ function is applied on the similarity map along the class dimension, thereby assigning each pixel to a particular class.
@@ 4 Experiments
4.1 Datasets and Evaluation Metric
We employ mean intersection-over-union (mIoU) to evaluate the performance, which is a standard metric in semantic segmentation.
4.2 Implementation Details
In our implementation, we utilize the Vicuna [5] as the large language model (LLM) for extracting tags. The visual encoder is based on the CLIP ViT-B/16 model. The input images are resized to a resolution of $224\times 224$ pixels, and the patch size used is $16\times 16$. We follow the modifications described in [19] to adapt the CLIP image encoder for our method. For the projector, we adopt the design proposed in [3], which consists of two gated convolution blocks, but we exclude the upsampling interpolations in our implementation. In our framework, only the lightweight projector is trainable, while all other modules are fixed. All our experiments are conducted on a single node equipped with 8 NVIDIA A100 GPUs, with each GPU having 80GB of memory. We train the model with a batch size of $4,096$ and a constant learning rate of 1e-3 for a total of $30,000$ iterations. Notably, the learning rate in our approach is not adjusted using any heuristic tricks (e.g., warmup or decay schedule), thereby reducing the number of hyperparameters involved in the training process. During inference, following the CLIP-based method TCL [3], we combine the segmentation results of our method and those of original CLIP through a weighted sum. Please refer to the Appendix for more details.